1) What is its unemployment rate?  


2) Now suppose 4,000 of the people looking for work get discouraged and give up their searches. What happens to the unemployment rate?  


3) Would you interpret this as good news for the economy of bad news? Explain.


1
Expert's answer
2021-12-05T18:58:11-0500

1.

The unemployment rate is defined as the percentage of the workforce that is unemployed.

2.

Unemployment rate = 10,000/(90,000+10,000) = 10%

If4,000 of the people looking for work get discouraged and give up their searches, the unemployment rate will decrease:

Unemployment rate = 6,000/(90,000+6,000)= 6.25%

3.

It is good, because less people are searching work, but is bad, because the total labor force decreases.